The difference between two numbers is 10. If the numbers are doubled, what is the difference between them?
Found 3 solutions by stanbon, solver91311, Mathtut:
Answer by stanbon(48558)

The difference between two numbers is 10. If the numbers are doubled, what is the difference between them?
----------------------------
If x-y = 10
then 2x -2y = 20
=======================
Cheers,
Stan H.

Answer by solver91311(12126)

One number is x and the other is y. The problem states that the difference is 10, so x-y=10. But if you double the numbers, you have to double the difference as well in order for the equation to remain a true statement.

2x-2y=20.


Answer by Mathtut(3670)

a-b=10.....eq 1
:
2a-2b=?....eq 2
:
2(a-b)=? distribution property
2(10)=20 substitute eq 1's value